OpenStudy (anonymous):

Using the 8 balls, numbered from 1 to 8 from Question 1,) show that it is possible to make 6,720 different 5-digit numbers, using each number only once, (hint: The 'box' method may help you with this.) (this step is if we DON'T need every 'thing; - if we only want to arrange some of them.)

OpenStudy (anonymous):

http://openstudy.com/users/tl9800#/updates/530324dae4b024fe26e31150 This was question 1

OpenStudy (anonymous):

I am not supposed to use factorial for this one.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

I think it has to be the box method

OpenStudy (anonymous):

well then # of selections for the first digit = n = 8 # of selections for the second digit = n-1 =7 . . . # of selections for the fifth digit = n-4 =4 # of ways = 8*7*6*5*4=6720

OpenStudy (anonymous):

I'm not sure what box-method is. I tried googling.

OpenStudy (anonymous):

Thank you! I am pretty sure this is it!